[н] [ он р он pH 4.7 7.7 x10-3 2.3 х10-8 3.6
Answer #1

DAN O pOH = 14-pH= 14-4.7 = 9.3 I Cart] = 10-pH = 10 -4.752x10-5M [OH]=10-20M: 10-3-3 = 5.01x10–10M @p#= -109 (m+)= -109.77.